# Who Are We Missing?: A Principled Approach to Characterizing the Underrepresented Population

### v2

@@ Appendix A Proof of Results in Section 3
(b) Variance of $\widehat{\tau}_{0}^{ipw}$ Estimator
$\widehat{\tau}_{0}$ $=$ $\frac{1}{n_{1}}\sum_{i=1}^{n}\left(\frac{S_{i}T_{i}Y_{i}}{\ell(\mathbf{X}_{i})e(\mathbf{X}_{i})}-\frac{S_{i}(1-T_{i})Y_{i}}{\ell(\mathbf{X}_{i})(1-e(\mathbf{X}_{i}))}\right)\left(\frac{\pi}{1-\pi}\right)$
and $\text{and }{\tau}_{0}^{\mathcal{D}_{n}}$ $=$ $\frac{1}{n_{1}}\sum_{i=1}^{n}\frac{S_{i}}{\ell(\mathbf{X}_{i})}\left(Y_{i}(1)-Y_{i}(0)\right)\left(\frac{\pi}{1-\pi}\right).$
@@
$\widehat{\tau}_{0}-{\tau}_{0}^{\mathcal{D}_{n}}$ $=$ $\frac{1}{n_{1}}\left(\frac{\pi}{1-\pi}\right)\sum_{i=1}^{n}\left(\frac{S_{i}}{\ell(\mathbf{X}_{i})}\right)\left(\left(\frac{Y_{i}T_{i}}{e(\mathbf{X}_{i})}-Y_{i}(1)\right)-\left(\frac{Y_{i}(1-T_{i})}{(1-e(\mathbf{X}_{i}))}-Y_{i}(0)\right)\right).$
Let $\psi_{i}=\left(\frac{S_{i}}{\ell(\mathbf{X}_{i})}\right)\left(\left(\frac{Y_{i}T_{i}}{e(\mathbf{X}_{i})}-Y_{i}(1)\right)-\left(\frac{Y_{i}(1-T_{i})}{(1-e(\mathbf{X}_{i}))}-Y_{i}(0)\right)\right)$ then,
$\widehat{\tau}_{0}-{\tau}_{0}^{\mathcal{D}_{n}}$ $=$ $\frac{1}{n_{1}}\left(\frac{\pi}{1-\pi}\right)\sum_{i=1}^{n}\psi_{i}.$
Here are interested in estimating $\mathbb{V}^{ipw}=Var(\sqrt{n_{1}}(\widehat{\tau}_{0}-{\tau}_{0}^{\mathcal{D}_{n}}))=n_{1}\mathbb{E}[\left(\widehat{\tau}_{0}-{\tau}_{0}^{\mathcal{D}_{n}}\right)^{2}]$.
$\mathbb{E}[\left(\widehat{\tau}_{0}-{\tau}_{0}^{\mathcal{D}_{n}}\right)^{2}\mid S_{i}=1]$ $=$ $\frac{1}{n_{1}^{2}}\left(\frac{\pi}{1-\pi}\right)^{2}\mathbb{E}\left[\left(\sum_{i}\psi_{i}\right)^{2}\right]$
$=$ $\frac{1}{n_{1}^{2}}\left(\frac{\pi}{1-\pi}\right)^{2}\sum_{i}\mathbb{E}\left[\psi_{i}^{2}\mid S_{i}=1\right]$
$+2\sum_{i}\sum_{j<i}\mathbb{E}\left[\psi_{i}\mid S_{i}=1\right]\mathbb{E}\left[\psi_{j}\mid S_{j}=1\right].$
As $\mathbb{E}\left[\psi_{i}\mid S_{i}=1\right]=0$ and $\psi_{i}\perp\psi_{j}$,
$\mathbb{E}[\left(\widehat{\tau}_{0}-{\tau}_{0}^{\mathcal{D}_{n}}\right)^{2}\mid S_{i}=1]$ $=$ $\frac{1}{n_{1}^{2}}\left(\frac{\pi}{1-\pi}\right)^{2}\sum_{i}\mathbb{E}\left[\psi_{i}^{2}\mid S_{i}=1\right]$
$\mathbb{E}[\left(\widehat{\tau}_{0}-{\tau}_{0}^{\mathcal{D}_{n}}\right)^{2}\mid S_{i}=1]$ $=$ $\frac{1}{n_{1}}\left(\frac{\pi}{1-\pi}\right)^{2}\mathbb{E}\left[\psi^{2}\mid S=1\right].$
Simplifying, $\psi$:
$\psi$ $=$ $\left(\frac{S}{\ell(\mathbf{X})}\right)\left[\Delta_{1}-\Delta_{0}\right]$
where,
$\Delta_{1}$ $=$ $\frac{YT}{e(X)}-Y(1)$
$\Delta_{0}$ $=$ $\frac{Y(1-T)}{1-e(X)}-Y(0).$
Thus,
$\psi^{2}$ $=$ $\left(\frac{S}{\ell(\mathbf{X})}\right)^{2}\left[\Delta_{1}^{2}+\Delta_{0}^{2}-2\Delta_{1}\Delta_{0}\right].$
$\mathbb{E}[\Delta_{1}^{2}\mid\mathbf{X},S=1]$ $=$ $\mathbb{E}[Y^{2}(1)\mid\mathbf{X},S=1]\left[\frac{1-e(\mathbf{X})}{e(\mathbf{X})}\right]$
$\mathbb{E}[\Delta_{0}^{2}\mid\mathbf{X},S=1]$ $=$ $\mathbb{E}[Y^{2}(0)\mid\mathbf{X},S=1]\left[\frac{e(\mathbf{X})}{1-e(\mathbf{X})}\right]$
$\mathbb{E}[\Delta_{1}\Delta_{0}\mid\mathbf{X},S=1]$ $=$ $-\mathbb{E}[Y(1)Y(0)\mid\mathbf{X},S=1]$
Assuming independence of potential outcomes,
$\mathbb{E}[\Delta_{1}\Delta_{0}\mid\mathbf{X},S=1]=-\mathbb{E}[Y(1)\mid\mathbf{X},S=1]\mathbb{E}[Y(0)\mid\mathbf{X},S=1].$
Putting these pieces together,
$\mathbb{E}[\psi^{2}\mid\mathbf{X},S=1]$ $=$ $\mathbb{E}[Y^{2}\mid T=1,\mathbf{X},S=1]\frac{1-e(\mathbf{X})}{e(\mathbf{X})}$
$+\mathbb{E}[Y^{2}\mid T=0,\mathbf{X},S=1]\frac{e(\mathbf{X})}{1-e(\mathbf{X})}$
$+2\mathbb{E}[Y\mid T=1,\mathbf{X},S=1]\mathbb{E}[Y\mid T=0,\mathbf{X},S=1]$
$=$ $\left(\mathbb{E}[Y^{2}\mid T=1,\mathbf{X},S=1]-\mathbb{E}^{2}[Y\mid T=1,\mathbf{X},S=1]\right)\frac{1-e(\mathbf{X})}{e(\mathbf{X})}$
$+\left(\mathbb{E}[Y^{2}\mid T=0,\mathbf{X},S=1]-\mathbb{E}^{2}[Y\mid T=0,\mathbf{X},S=1]\right)\frac{e(\mathbf{X})}{1-e(\mathbf{X})}$
$+\mathbb{E}^{2}[Y\mid T=1,\mathbf{X},S=1]\frac{1-e(\mathbf{X})}{e(\mathbf{X})}$
$+\mathbb{E}^{2}[Y\mid T=0,\mathbf{X},S=1]\frac{e(\mathbf{X})}{1-e(\mathbf{X})}$
$+2\mathbb{E}[Y\mid T=1,\mathbf{X},S=1]\left(\frac{1-e(\mathbf{X})}{e(\mathbf{X})}\right)^{1/2}\mathbb{E}[Y\mid T=0,\mathbf{X},S=1]\left(\frac{e(\mathbf{X})}{1-e(\mathbf{X})}\right)^{1/2}.$
Thus,
$\mathbb{V}^{ipw}=\left(\frac{\pi}{1-\pi}\right)^{2}\mathbb{E}_{\mathbf{X}\mid S=1}\left[\frac{\sigma^{2}(\mathbf{X},1)\frac{1-e(\mathbf{X})}{e(\mathbf{X})}+\sigma^{2}(\mathbf{X},0)\frac{e(\mathbf{X})}{1-e(\mathbf{X})}+\left(\mu(\mathbf{X},1)\left(\frac{1-e(\mathbf{X})}{e(\mathbf{X})}\right)^{1/2}+\mu(\mathbf{X},0)\left(\frac{e(\mathbf{X})}{1-e(\mathbf{X})}\right)^{1/2}\right)^{2}}{(\ell(\mathbf{X}))^{2}}\right]$

### v4

@@ Appendix A Proof of Results in Section 3
(b) Variance of TATE IPW Estimate $\widehat{\tau}_{0}$
$\widehat{\tau}_{0}=\frac{1}{n}\sum_{i=1}^{n}\psi_{i}\text{ where }\psi_{i}=\left(\frac{S_{i}T_{i}Y_{i}}{e(X_{i})\pi(X_{i})}-\frac{S_{i}(1-T_{i})Y_{i}}{(1-e(X_{i}))\pi(X_{i})}\right)\frac{(1-\pi(X_{i}))}{(1-\pi)}.$
Now, consider
$B=\widehat{\tau}_{0}-{\tau}_{0}=\frac{1}{n}\sum_{i=1}^{n}(\psi_{i}-{\tau}_{0})=\frac{1}{n}\sum_{i=1}^{n}(\psi_{i}-{\tau}_{0}(X_{i})+{\tau}_{0}(X_{i})-{\tau}_{0})$
where ${\tau}_{0}(X_{i})=\frac{1-\pi(X_{i})}{1-\pi}\left(\mu(X_{i},1)-\mu(X_{i},0)\right)$ Let $\sigma^{2}(x,t)$ be the variance of $Y$ conditional on $X=x$, $T=t$ and $S=1$.
$Var(\widehat{\tau}_{0}-{\tau}_{0})=\mathbb{E}[B^{2}]=\frac{1}{n}Var(\psi-{\tau}_{0}(X)+{\tau}_{0}(X)-{\tau}_{0})$
$Var(\psi-{\tau}_{0}(X)+{\tau}_{0}(X)-{\tau}_{0})=\mathbb{E}_{X}[\mathbb{E}[(\psi-{\tau}_{0}(X)+{\tau}_{0}(X)-{\tau}_{0})^{2}\mid X]]$
$=\mathbb{E}_{X}[\mathbb{E}[(\psi-{\tau}_{0}(X))^{2}\mid X]]+\mathbb{E}_{X}[\mathbb{E}[({\tau}_{0}(X)-{\tau}_{0})^{2}\mid X]]$
Now, first let’s consider $\mathbb{E}_{X}[\mathbb{E}[(\psi-{\tau}_{0}(X))^{2}\mid X]]$:
$=\mathbb{E}_{X}\left[\mathbb{E}\left[\left(\frac{1-\pi(x)}{1-\pi}\right)^{2}\left(\frac{STY}{\pi(x)e(X)}-\mu(X,1)-\frac{S(1-T)Y}{\pi(x)(1-e(X))}+\mu(X,0)\right)^{2}\mid X\right]\right]$
$=\mathbb{E}_{X}\left[\left(\frac{\sigma^{2}(X,1)}{e(X)\pi(X)}+\frac{\sigma^{2}(X,0)}{(1-e(X))\pi(X)}\right)\left(\frac{(1-\pi(X))}{(1-\pi)}\right)^{2}\right]$
$=\frac{\pi}{(1-\pi)^{2}}\mathbb{E}_{X|S=1}\left[\left(\frac{\sigma^{2}(X,1)}{e(X)}+\frac{\sigma^{2}(X,0)}{(1-e(X))}\right)\left(\frac{(1-\pi(X))}{\pi(X)}\right)^{2}\right]$
Focusing on $\mathbb{E}_{X}[\mathbb{E}[({\tau}_{0}(X)-{\tau}_{0})^{2}\mid X]]$:
$\mathbb{E}_{X}[\mathbb{E}[({\tau}_{0}(X)-{\tau}_{0})^{2}\mid X]]=\mathbb{E}_{X}\left[\left(\frac{1-\pi(X)}{1-\pi}(\mu(X,1)-\mu(X,0))-{\tau}_{0}\right)^{2}\right]$
$=\mathbb{E}_{X|S=1}\left[\frac{\pi}{\pi(X)}\left({\tau}_{0}(X)-{\tau}_{0}\right)^{2}\right]$
@@
$Var(\widehat{\tau}_{0}-{\tau}_{0})$ $=$ $\frac{\pi}{(1-\pi)^{2}}\mathbb{E}_{X|S=1}\left[\left(\frac{\sigma^{2}(X,1)}{e(X)}+\frac{\sigma^{2}(X,0)}{(1-e(X))}\right)\left(\frac{(1-\pi(X))}{\pi(X)}\right)^{2}\right]$
$+\mathbb{E}_{X|S=1}\left[\frac{\pi}{\pi(X)}\left({\tau}_{0}(X)-{\tau}_{0}\right)^{2}\right]$
@@ Appendix B ROOT Algorithm
Appendix C Consistency of Tree-based Optimization
In this section, we discuss the convergence of random tree sampling-based optimization such as ROOT. ROOT samples random trees that favor splits on features correlated with the outcome i.e. the probability of splitting on a feature is equal to the correlation between the feature and the outcome. However, here, we show proof for a pure random tree with $Q$ leaves (PRTQ) sampling approach that splits uniformly over each feature. This is done for mathematical simplicity, and the PRTQ sampling approach is a special case of ROOT. Furthermore, trees sampled via a pure random tree sampling approach are independent and identically distributed.
Similar to the literature on tree (Xin et al.,, 2022; Lin et al.,, 2020) we consider $\mathbf{X}_{n\times p}$ to be a $n\times p$-dimensional matrix of pretreatment covariates with binary entries where row correspond to different units and columns correspond to different features. Note that you can convert any finite matrix of pre-treatment covariates into a matrix with binary entries (Lin et al.,, 2020). Now, we describe a pure random tree sampling approach to sample a tree with $Q$ leaves (Breiman,, 2000):
  1. 1.
    At step 00, we start with a root node
  2. 2.
    At any $(q-1)$-th step, there are $q$ existing leaf nodes
    1. (a)
      Randomly choose one of them to split and grow the tree
    2. (b)
      Randomly choose one of the $p$ features to split on
  3. 3.
    At $(Q-1)$-th step, the constructed tree has $Q$ leaves
    1. (a)
      For each leave, randomly choose the label from $\{0,1\}$.
[…]
Let $w^{Q}_{*}\in\arg\min_{w^{Q}}\mathcal{L}(w^{Q},\mathcal{D}_{n})$ denote the optimal tree with $Q$-leaves that minimizes the objective function $\mathcal{L}$. Let $\mathcal{W}^{Q}_{K}:=\{w_{1}^{Q},\dots,w_{K}^{Q}\}$ denote a set of $K$ independently and identically sampled trees. Then, we are interested in understanding if $w^{Q}_{*}\in\mathcal{W}^{Q}_{K}$ with high probability.
For any tree $w^{Q}_{k}\in\mathcal{W}^{Q}_{K}$,
$P(w^{Q}_{k}=w^{Q}_{*})\geq P(\{o^{k}_{1},\dots,o^{k}_{Q-1}\}=\{o^{*}_{1},\dots,o^{*}_{Q-1}\})P(\{l^{k}_{1},\dots,l^{k}_{Q}\}=\{l^{*}_{1},\dots,l^{*}_{Q}\}).$
At q-th step, $P(o^{k}_{q}=o^{*}_{q}\mid\{o^{k}_{1},\dots,o^{k}_{q-1}\}=\{o^{*}_{1},\dots,o^{*}_{q-1}\})=\frac{1}{q-1}\frac{1}{p}$. Thus, $P(\{o^{k}_{1},\dots,o^{k}_{Q-1}\}=\{o^{*}_{1},\dots,o^{*}_{Q-1}\})=\frac{p}{Q-1!}\left(\frac{1}{p}\right)^{Q}$. Further, $P(\{l^{k}_{1},\dots,l^{k}_{Q}\}=\{l^{*}_{1},\dots,l^{*}_{Q}\})=\frac{1}{2^{Q}}$. Thus $P(w^{Q}_{k}=w^{Q}_{*})\geq\frac{p}{Q-1!}\left(\frac{1}{2p}\right)^{Q}$ and $P(w^{Q}_{k}\neq w^{Q}_{*})\leq 1-\frac{p}{Q-1!}\left(\frac{1}{2p}\right)^{Q}.$ Thus,
$P(w^{Q}_{*}\notin\mathcal{W}^{Q}_{K})\leq\left(1-\frac{p}{Q-1!}\left(\frac{1}{2p}\right)^{Q}\right)^{K}.$
Thus, as number of sampled trees, $K$, goes to $\infty$, $P(w^{Q}_{*}\notin\mathcal{W}^{Q}_{K})$ goes to 00.
Note that these bounds are very loose and for a weaker tree sampling algorithm than ROOT. Further, theoretical work will focus on exploring a tighter bound and a better convergence rate for ROOT.
Appendix D Synthetic Data Experiments
Now, we study the performance of our proposed approach via synthetic data experiments. In particular, we are interested in understanding the extent to which our approach (i) produces interpretable characterizations of underrepresented populations, (ii) reduces the variance of the WTATE relative to the variance of the TATE, and (iii) handles high-dimensional and non-linear data in four data-generative processes (DGPs) with different features.
D.1 Data Generative Process
These four synthetic DGPs are designed to be akin to different real-world scenarios and illustrate important features of our approach. We call them (i) Community, (ii) Box, (iii) Clone, and (iv) High-dimensional DGPs due to their respective properties.
Community DGP.
$P(C_{i}=A)=\frac{3}{4}\textrm{ and }P(C_{i}=B)=\frac{1}{4}$
Box DGP.
$X_{i,0},X_{i,1},\ldots,X_{i,100}\sim\mathcal{U}(0,1).$
$Y_{i}(1)=Y_{i}(0)+\log(Y_{i}(0)+1).$
Note that, here, covariates $X_{i,0}$ and $X_{i,1}$ are affecting both sample selection $S$ and the treatment effect $Y_{i}(1)-Y_{i}(0)$. In summary, the Box DGP model involves high-dimensional covariates, non-linear response functions, and varying trial participation probabilities based on covariate characteristics.
Clone DGP.
We design this DGP to mimic the data distributions in the case study of our interest (presented in Section 6.2). We provide a detailed description of data in Section 6.2, however, here, it is important to know that $Y$ is a binary outcome and $\mathbf{X}$ is a 10-dimensional vector. To generate the synthetic data that mimics our observed data, we need to design the joint distribution $P(Y(1),Y(0),\mathbf{X},S,T)$. Leveraging assumption A.2. and A.3., we rewrite the joint as $P(Y(1),Y(0),\mathbf{X},S,T)=P(T\mid S,\mathbf{X})P(S\mid\mathbf{X})P(Y(1),Y(0),\mathbf{X})$. We learn the following models using the observed MOUD data: (i) $P(Y(1)=1\mid\mathbf{X},S=1)$ (ii) $P(Y(0)=1\mid\mathbf{X},S=1]$, (iii) $P(T=1\mid\mathbf{X},S=1)$ and (iv) $P(S=1\mid\mathbf{X})$. We use gradient-boosting trees to learn the two potential outcome models and cross-validated logistic regression to learn the propensity and selection scores. We use the potential outcome models to impute the missing potential outcomes in the observed data and then learn the joint distribution $P(Y(1),Y(0),\mathbf{X})$. And with a slight abuse of notations, to sample a synthetic data with $n$ units, we first sample $\{Y_{i}(1),Y_{i}(0),\mathbf{X}_{i}\}_{i=1}^{n}\overset{iid}{\sim}P(Y(1),Y(0),\mathbf{X})$. Next, we sample $\{S_{i}\}$ and then, $\{T_{i}\}$ for units with $S_{i}=1$. This data generative procedure guarantees that the empirical distributions of the observed data and the synthetic data are similar to each other. Further, the trained models inform the ground-truth effect modifiers as well as sample selectors for the synthetic data. Figure 4 shows the relative importance of each feature in determining the selection probability (y-axis) and treatment effect (x-axis).
Figure 4: Scatter plot of Clone DGP relative feature importance for sample selection and treatment effect.
High-dimensional DGP.
Similar to the Box DGP, here too we consider a high-dimensional setup with 100 iid covariates per individual: $\{X_{i,0}\dots X_{i,100}\}\overset{iid}{\sim}\mathcal{N}(0,1)$. Unlike the Box DGP, all 100 covariates contribute towards an individual’s likelihood of participating in the trial: $P(S_{i}=1\mid\mathbf{X}_{i})=\mathrm{expit}(\alpha_{0}^{T}\mathbf{X}_{i})$. The control potential outcome and the treatment effect are linear functions of covariates: $Y_{i}(0)=\beta_{0}^{T}\mathbf{X}_{i}+\epsilon_{i}$, and $Y_{i}(1)=Y_{i}(0)+\beta_{1}^{T}\mathbf{X}_{i}+\epsilon_{i}$ where $\epsilon_{i}\sim\mathcal{N}(0,1)$. Here, the parameters $\alpha_{0}$, $\beta_{0}$, and $\beta_{1}$ are also drawn from a standard normal distribution. The analysis challenge in this DGP is that the features affecting the sample selection and treatment effect are high-dimensional. Thus there is no sparse representation of the underrepresented group.
D.2 Analysis
We first examine the extent to which our approach reduces the variance of the WTATE estimates compared to the TATE estimates. Table 2 shows the standard errors for each of the four approaches as well as the estimates on the unpruned (or unrefined) target. The precision measured as the standard error is a function of study data as well as the estimator. We use the same IPW WTATE estimator across all 4 approaches which is an unbiased estimator of treatment effect. Focusing on the standard errors (precision), we find that for Community, Box, and Clone DGPs, the ROOT approach produces the most precise estimates. For High-dimensional DGP, however, the Linear-w approach tends to outperform other approaches. This is primarily because all the relationships in this DGP are linear and thus $w(\mathbf{X})=\gamma^{T}\mathbf{X}$ is the “correct” specification.
Thus, our experimental results indicate that our approach improves the precision of the TATE estimate without compromising the bias. Further, given the nonlinearity of the Box DGP, and the high-dimensional nature of both the Box and High-dimensional DGPs, our results also support the ability of our approach to handle a complex and high-dimensional dataset.
Original
Target
Estimate
ROOT 1-Tree Linear Indicator
Community DGP
   Std. Error 3.073 0.449 0.985 2.995 0.770
Box DGP
   Std. Error 0.910 0.324 0.409 0.910 0.763
Clone DGP
   Std. Error 4.655 4.520 4.655 4.655 4.551
High-dim. DGP
   Std. Error 0.191 0.162 0.169 0.127 0.170
Table 2: Table of Synthetic Data Analysis Results. Precision (measured as standard error) for the four DGPs described in Section D.1, and four approaches described in Section 5. ROOT is our proposed approach while the other three are baseline comparison methods.
Next, we discuss how one can use the ROOT approach to produce interpretable characteristics of the underrepresented groups. In Figure 5, we shows trees characterizing the insufficiently represented populations for the four DGPs. Further, Figure 6 shows the data distribution color-coded based on the predicted values of $w(\mathbf{X})$ using ROOT for each unit in the dataset across the two most important features for each of the DGPs that are both outcome modifiers as well as the sample selectors. Our results shown in these figures indicate that ROOT is not only correctly able to identify the underrepresented groups or communities but also produces a tree that can be used to communicate their characteristics easily.
Unlike the results for the Community, Box, and Clone DGPs, in the case of the High-dimensional DGP, the decision boundary in the two-dimensional projection is not sharp. This is because in this scenario all of the 100 covariates are moderators and relate to trial participation and no two-dimensional slicing of the covariate space can perfectly show the decision boundary.
Further, for the high-dimensional DGP, it is worth noting that ROOT is only the second-best approach in terms of precision even this DGP is designed to be challenging for ROOT. This is because the default hyperparameters of ROOT are chosen to find subgroups with sparse representations (i.e., the ones described via shallow trees) for the sake of interpretability. Due to a large number of moderators being almost equally important in the construction of this DGP, the shallow representation yields a less-than-optimal representation of the underrepresented population. This is because of the large number of moderators being almost equally important and thus there is no shallow tree that can accurately characterize the underrepresented groups.
(a) Community DGP
(b) Box DGP
(c) Clone DGP
(d) High-dim. DGP
Figure 5: Tree characterizing indicating the study population in Blue and the underrepresented subgroups in Orange for all four DGPs
Refer to caption Refer to caption Refer to caption
(a) Community DGP (b) Box DGP (c) High-dim. DGP
Appendix E MOUD Data Descriptive Statistics
We consider methadone as $T=1$ and buprenorphine as $T=0$. For the outcome, $Y=1$ indicates relapse by 24 weeks after treatment assignment. We only consider the pre-treatment covariates observed in both the trial and target data. Our set of pre-treatment covariates includes the participants’ age, race and biological sex along with substance use history (amphetamine, benzodiazepines, cannabis and IV drug use) assessed at the initiation of MOUD treatment. We provide Table 3 for a listing of all of the covariates included.
TEDS-A START
S=0 S=1
Sample Size 740,015 1,271
Covariates
   Male (%) 58% 68%
   Age (IQR) [27y,47y] [27y,46y]
   History of Benzo Use (%) 6% 17%
   History of Amphetamine Use (%) 7% 13%
   History of IV Drug Use (%) 54% 70%
   History of Cannabis Use (%) 11% 29%
   Race
    White (%) 65% 68%
    Black (%) 15% 9%
    Hispanic (%) 14% 16%
    Other Race (%) 5% 8%
Treatment
   Methadone (%) - 42%
   Buprenorphine (%) - 58%
Outcome
   Relapse in 24w (%) - 71%
Table 3: The distributions of demographic and clinical pre-treatment covariates, treatments, and outcomes across the START trial and TEDS-A datasets.
Appendix F Discussion on Assumptions for MOUD Case Study
Figure 7: Violin-plot showing the distribution of the normalized selection odds $\ell(\mathbf{x})/\ell$ for the units in trial and target samples.
  • (A1) This assumption can be assessed through examination of the data; Figure 4 depicts the overlap between the distributions of the target population and trial cohort on normalized conditional selection odds (i.e., $\frac{\ell(x)}{\ell}=\frac{1-\pi}{\pi}\frac{\pi(x)}{1-\pi(x)}$). We observe that there is significant overlap of the estimated selection odds between the trial and target samples. Further, none of the units have selection probabilities and thus selection odds equal to 0. Thus, structural positivity assumption (A1) appears to hold in this data, i.e., $P(S=1\mid X=x)\neq 0$ for all units in the target sample.
  • (A2) In our study, we focus on transporting the intent-to-treat effect using the START trial data. Given the randomized design of the trial Saxon et al., (2013), unconfounded treatment assignment (A2) holds.
  • (A3) A3 assumes the absence of an unobserved factor, $U$, affecting both the outcome as well as the trial selection probability once the set of observed covariates $x$ are adjusted for. This assumption is untestable. Based on domain knowledge, accounting for $x$ may be sufficient to ensure the exchangeability of the potential outcomes across $S=0$ and $S=1$ (Rudolph et al.,, 2021, 2022; Nunes Jr et al.,, 2021; Rudolph et al., 2023a, ). However, certain potentially important variables such as homelessness, parole and probation status, medication preference are unobserved across datasets and could potentially lead to violation of A3 (Nunes Jr et al.,, 2021).
  • (A4) Stable unit treatment value assumption (SUTVA) has two parts: (i) there is no interference and (ii) there is a single version of each of the treatments across all units. The first part may be violated because 1) individuals are enrolled across 8 opioid treatment programs, and 2) within each treatment program, individuals may be part of common social networks. The same limitations regarding patients nested within program and with overlapping social networks exist in the TEDS-A data. How to best address these complex dependencies is an area for future work. For now we assume any such SUTVA violation is negligible. Further, the second part may be challenged because the MOUD doses are not standardized across treatment programs. However, our work focuses on estimating intent to treatment effect and thus, we assume that this part of SUTVA holds as well.